A typical family living on Rabbit Island consumes only apple juice, bananas, and cloth. Price in the base year are $4 a litre for apple juice, $3 a kilogram for bananas, and $5 a metre for cloth. In the base year the typical family spends $40 on apple juice, $45 on bananas, and $25 on cloth. In the current year, apple juice cost $3 a litre, banans cost $4 a kilogram, and cloth costs $7 a metre. Calculate the CPI on Rabbit Island in the current year and inflation rate between the base year and the current year.
1
Expert's answer
2019-01-21T10:44:10-0500

Price in the base year are $4 a litre for apple juice, $3 a kilogram for bananas, and $5 a metre for cloth. In the base year the typical family spends $40 on apple juice, $45 on bananas, and $25 on cloth, so quantities consumed are: 40/4 = 10 litres of apple juice, 45/3 = 15 kilogram of bananas, 25/5 = 5 metres of cloth.

In the current year, apple juice cost $3 a litre, bananas cost $4 a kilogram, and cloth costs $7 a metre.

The CPI on Rabbit Island in the current year is:

CPI = ($3*10 + $4*15 + $7*5)/(40 + 45 + 25)*100 = 113.6.

Inflation rate between the base year and the current year is 13.6%.